OpenStudy (anonymous):

ln(x^18 sqrt(y^14/z^19))

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

to expand using the laws of logarithms

OpenStudy (anonymous):

some of the laws are: log(A) + log(B) = log(AB) log(A) - log(B) = log(A/B) log(A^n) = nlogA the bases of the log have to be the same for this to work

OpenStudy (anonymous):

is ln and log the same thing

OpenStudy (anonymous):

pretty much. ln is log base e. the log button on your calculator is log base 10 while ln is log base e.

OpenStudy (anonymous):

the answer i get is: 18lnx + 7lny - (19/2)lnz

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

lmk if you would like to know how i got the answer after trying to reproduce it

OpenStudy (anonymous):

yeah can you show me i got the first part but i made a mistake on the z